# FAQNAS: FLOPs-aware Hybrid Quantum Neural Architecture Search using Genetic Algorithm

### v1

@@ I Introduction
The rapid advancement of quantum computing has spurred significant interest in Quantum Machine Learning (QML) as a paradigm that may offer computational advantages over classical methods in the near term [1]. Hybrid Quantum Neural Networks (HQNNs), which integrate parameterized quantum circuits (PQCs) with classical neural layers (Fig. 1), have emerged as promising candidates for leveraging noisy intermediate-scale quantum (NISQ) devices [2]. These architectures combine the representational power of quantum states with the scalability of classical models, making them attractive for practical learning tasks. However, the design of effective HQNN architectures remains a non-trivial challenge. Unlike classical deep learning, where architecture search has been extensively studied [3, 4], architecture design in HQNNs is still relatively underexplored. Existing efforts in quantum architecture search have primarily targeted performance optimization, either by customizing circuit designs for specific devices or by developing strategies to mitigate hardware noise and scalability constraints, with the aim of achieving higher accuracy [5, 6, 7]. As most of these approaches are performance-oriented, our methodology fills this gap by giving equal importance to predictive accuracy and computational resource efficiency, thereby enabling a more balanced design.
As HQNNs continue to evolve, a central question emerges: how can we design architectures that deliver competitive performance while keeping computational demands manageable? In practice, model selection for HQNNs involves conflicting objectives. On one hand, increasing the number of trainable parameters or circuit depth can improve expressivity and classification performance[8]. On the other hand, deeper circuits and larger parameter spaces are costly: they exacerbate barren plateaus in the optimization landscape and introduce significant computational overhead[9, 10, 11]. This overhead is particularly critical in the NISQ era, where most HQNNs are trained and benchmarked on classical simulators. In such settings, floating-point operations (FLOPs) directly determine runtime feasibility, memory footprint, and scalability[12]. Even when real quantum hardware is used for circuit execution, the hybrid training loop, including gradient estimation, optimization, and loss evaluation, remains largely classical and FLOPs-intensive. Thus, FLOPs provide a practical and unifying measure of computational cost in HQNNs.
@@
Figure 1: An illustration Hybrid Quantum Neural Network Architecture. Typically Parametrized Quantum circuits are sandwiched between classical neuron layers with classical optimization loop.
@@ II-B Datasets
To evaluate the effectiveness and generalizability of the proposed framework, we employed three widely used benchmark datasets that vary in complexity and dimensionality:
@@ II-B4 Wine Dataset
The Wine dataset [15] contains $178$ samples of three wine cultivars. Each sample is described by $13$ continuous physicochemical features, such as alcohol content, flavanoids, and color intensity. This dataset include $3$ classes, corresponding to the wine cultivars, and represent a medium-sized classification challenge and enables assessment of the framework on structured tabular data.
@@ II-C HQNN Architecture
    Classical Pre-processing: The classical front-end consists of a fully connected (linear) layer that prepares raw input data for quantum processing. Given an input sample represented as a feature vector of dimensionality $d$, this layer projects it onto an $n_{\text{qubits}}$ dimensional feature space through multiplication with a weight matrix of size $d\times n_{\text{qubits}}$). The resulting vector of length $n_{\text{qubits}}$ serves as the embedding that is fed into the quantum layer. This linear transformation serves a dual purpose: dimensionality reduction for high-dimensional inputs and feature expansion for low-dimensional inputs, thereby aligning the data representation with the quantum circuit size.
@@
    Quantum Layers: The quantum layer forms the core of the hybrid architecture and is implemented using pennylane as a QNode wrapped as a PyTorch layer [^2] https://pennylane.ai/qml/demos/tutorial_qnn_module_torch, enabling seamless integration with classical optimization routines. The input vector from the front-end is embedded into the quantum state using one of two encoding strategies: angle embedding or amplitude embedding (Eq.1 and Eq.2, respectively), as selected by the Genetic Algorithm (GA). The PQC then applies up to a configurable number of layers of parameterized single-qubit rotations. For each qubit and each layer, the GA selects the gate type from the set $\{R_{X}(\theta),R_{Y}(\theta),R_{Z}(\theta)\}$. The rotation angles are trainable parameters organized in a tensor of shape $(\text{max\_layers},n_{\text{qubits}})$.
@@
    Classical Post Processing: The classical post processing module is responsible for mapping the quantum-transformed feature vector to the final prediction space. A fully connected layer projects the $n_{\text{qubits}}$-dimensional output into a vector of logits corresponding to the number of classes in the target dataset. A log-softmax activation is applied to convert the logits into log-probabilities, which are then used for training and evaluation through a negative log-likelihood (NLL) loss.
@@ II-D2 FAQNAS Work Flow
For a given $\mathbf{c}$, we build a three-stage HQNN similar to that shown in Fig. 1 and explained in Section II-C: (i) a classical linear (preprocessing) layer that flattens the input (e.g., $28{\times}28$ for MNIST) and projects to $\mathbb{R}^{n_{\text{qubits}}}$, (ii) a quantum module with the chosen feature map (Angle Embedding or Amplitude Embedding; for the latter we zero-pad and normalize to length $2^{n_{\text{qubits}}}$), $L$ layers of per-wire single-qubit rotations with the specified gate types, and $n_{\text{ent}}$ entangling layers under the selected topology; measurements are $Z$-expectations per qubit, and (iii) a linear classical (postprocessing) layer mapping $\mathbb{R}^{n_{\text{qubits}}}\to\mathbb{R}^{n}$ (where $n$ is number of classes in the dataset) followed by a negative log-likelihood (NLL) loss.
@@
To make FLOPs comparable across candidates, we measure one-sample forward+backward cost with PyTorch profiler [^3] https://docs.pytorch.org/docs/stable/profiler.html. For each genome, (i) We run the full hybrid model to obtain total number FLOPs ($F^{\text{Total}}$) and (ii) run a classical-only baseline (pre and post processing classical layers) to obtain the FLOPs consumption of classical layers $F^{\text{Classical}}$. The quantum contribution is $F^{\text{Q}}=\max\{F^{\text{Total}}-F^{\text{Classical}},0\}$. This isolates the incremental cost of the quantum layer while holding classical work constant except for the width of hidden layer coupling to $n_{\text{qubits}}$.
@@
  • We initialize $P$ genomes by uniform sampling of $\mathcal{S}$ and evaluate their objectives.
@@ III-F Comparative Analysis
A comparative analysis across datasets highlights several consistent trends. Accuracy generally increases with FLOPs but saturates after a dataset-dependent threshold. Large-scale datasets such as MNIST require higher FLOPs budgets to reach strong performance, while moderate datasets like Digits achieve near-maximum accuracy with relatively few FLOPs. Smaller and simpler datasets, including Wine, Breast Cancer, and Iris, frequently admit efficient solutions that achieve high accuracy at surprisingly low FLOPs, though with more variability in performance among candidate architectures. In all cases, improvements in accuracy are primarily driven by quantum FLOPs, since classical FLOPs remain largely fixed due to the constrained design of classical layers.
Overall, these results demonstrate that FLOPs-aware HQNN search consistently identifies architectures that balance accuracy and computational efficiency. The Pareto fronts across datasets confirm that strong performance can be obtained without over-provisioning FLOPs, with dataset complexity dictating the threshold at which additional FLOPs cease to provide accuracy gains. The findings underscore the importance of incorporating FLOPs-awareness into HQNN search, as it not only enables scalable optimization for large datasets but also prevents wasted computational resources for simpler ones.

### v2

@@ I Introduction
The rapid advancement of quantum computing has spurred significant interest in Quantum Machine Learning (QML) as a paradigm that may offer computational advantages over classical methods in the near term [23, 15]. Hybrid Quantum Neural Networks (HQNNs), which integrate parameterized quantum circuits (PQCs) with classical neural layers (Fig. 1), have emerged as promising candidates for leveraging noisy intermediate-scale quantum (NISQ) devices [5]. These architectures combine the representational power of quantum states with the scalability of classical models, making them attractive for practical learning tasks [6]. However, the design of effective HQNN architectures remains a non-trivial challenge. Unlike classical deep learning, where architecture search has been extensively studied [1, 22], architecture design in HQNNs is still relatively underexplored. Existing efforts in quantum architecture search have primarily targeted performance optimization, either by customizing circuit designs for specific devices or by developing strategies to mitigate hardware noise and scalability constraints, with the aim of achieving higher accuracy [2, 18, 17, 13, 12, 3]. As most of these approaches are performance-oriented, our methodology fills this gap by giving equal importance to predictive accuracy and computational resource efficiency, thereby enabling a more balanced design.
As HQNNs continue to evolve, a central question emerges: how can we design architectures that deliver competitive performance while keeping computational demands manageable? In practice, model selection for HQNNs involves conflicting objectives. On one hand, increasing the number of trainable parameters or circuit depth can improve expressivity and classification performance[8]. On the other hand, deeper circuits and larger parameter spaces are costly as they exacerbate barren plateaus in the optimization landscape and introduce significant computational overhead[20, 7, 14, 9, 11]. This overhead is particularly critical in the NISQ era, where most HQNNs are trained and benchmarked on classical simulators. In such settings, floating-point operations (FLOPs) directly determine runtime feasibility, memory footprint, and scalability[10]. Even when real quantum hardware is used for circuit execution, the hybrid training loop, including gradient estimation, optimization, and loss evaluation, remains largely classical and FLOPs-intensive. Thus, FLOPs provide a practical and unifying measure of computational cost in HQNNs.
@@
Figure 1: An illustration of a Hybrid Quantum Neural Network architecture. Typically Parametrized Quantum Circuits are sandwiched between classical neuron layers with classical optimization loop.
@@ II-B Datasets
To evaluate the effectiveness and generalizability of the proposed framework, we employed five widely used benchmark datasets that vary in complexity and dimensionality:
@@ II-B4 Wine Dataset
The Wine dataset [21] contains $178$ samples of three wine cultivars. Each sample is described by $13$ continuous physicochemical features, such as alcohol content, flavanoids, and color intensity. This dataset includes $3$ classes, corresponding to the wine cultivars, and represent a medium-sized classification challenge and enables assessment of the framework on structured tabular data.
@@ II-C HQNN Architecture
    Classical Pre-processing: The classical front-end consists of a fully connected (linear) layer that prepares raw input data for quantum processing. Given an input sample represented as a feature vector of dimensionality $d$, this layer projects it onto an $n_{\text{qubits}}$-dimensional feature space through multiplication with a weight matrix of size $d\times n_{\text{qubits}}$. The resulting vector of length $n_{\text{qubits}}$ serves as the embedding that is fed into the quantum layer. This linear transformation serves a dual purpose: dimensionality reduction for high-dimensional inputs and feature expansion for low-dimensional inputs, thereby aligning the data representation with the quantum circuit size.
@@
    Quantum Layers: The quantum layer forms the core of the hybrid architecture and is implemented using PennyLane as a QNode wrapped as a PyTorch layer [^2] https://pennylane.ai/qml/demos/tutorial_qnn_module_torch, enabling seamless integration with classical optimization routines. The input vector from the front-end is embedded into the quantum state using one of two encoding strategies: angle embedding or amplitude embedding (Eq. (1) and Eq. (2), respectively), as selected by the Genetic Algorithm (GA). The PQC then applies up to a configurable number of layers of parameterized single-qubit rotations. For each qubit and each layer, the GA selects the gate type from the set $\{R_{X}(\theta),R_{Y}(\theta),R_{Z}(\theta)\}$. The rotation angles are trainable parameters organized in a tensor of shape $(\text{max\_layers},n_{\text{qubits}})$.
@@
    Classical Post-Processing: The classical post-processing module is responsible for mapping the quantum-transformed feature vector to the final prediction space. A fully connected layer projects the $n_{\text{qubits}}$-dimensional output into a vector of logits corresponding to the number of classes in the target dataset. A log-softmax activation is applied to convert the logits into log-probabilities, which are then used for training and evaluation through a negative log-likelihood (NLL) loss.
@@ II-D2 FAQNAS Workflow
For a given $\mathbf{c}$, we build a three-stage HQNN similar to that shown in Fig. 1 and explained in Section II-C: (i) a classical linear (pre-processing) layer that flattens the input (e.g., $28{\times}28$ for MNIST) and projects to $\mathbb{R}^{n_{\text{qubits}}}$, (ii) a quantum module with the chosen feature map (Angle Embedding or Amplitude Embedding; for the latter we zero-pad and normalize to length $2^{n_{\text{qubits}}}$), $L$ layers of per-wire single-qubit rotations with the specified gate types, and $n_{\text{ent}}$ entangling layers under the selected topology; measurements are $Z$-expectations per qubit, and (iii) a linear classical (postprocessing) layer mapping $\mathbb{R}^{n_{\text{qubits}}}\to\mathbb{R}^{n}$ (where $n$ is number of classes in the dataset) followed by a negative log-likelihood (NLL) loss.
@@
To make FLOPs comparable across candidates, we measure one-sample forward+backward cost with PyTorch profiler [^3] https://docs.pytorch.org/docs/stable/profiler.html. For each genome, (i) We run the full hybrid model to obtain total number FLOPs ($F^{\text{Total}}$) and (ii) run a classical-only baseline (pre and post-processing classical layers) to obtain the FLOPs consumption of classical layers $F^{\text{Classical}}$. The quantum contribution is $F^{\text{Q}}=\max\{F^{\text{Total}}-F^{\text{Classical}},0\}$. This isolates the incremental cost of the quantum layer while holding classical work constant except for the width of hidden layer coupling to $n_{\text{qubits}}$.
@@
  • We initialize $P$ genomes by uniform sampling of search space $\mathcal{S}$ and evaluate their objectives.
@@ III-F Comparative Analysis
A comparative analysis across datasets highlights several consistent trends. Accuracy generally increases with FLOPs but saturates after a dataset-dependent threshold. Large-scale datasets such as MNIST require higher FLOPs budgets to reach strong performance, while moderate datasets like Digits achieve near-maximum accuracy with relatively few FLOPs. Smaller and simpler datasets, including Wine, Breast Cancer, and Iris, frequently yields efficient solutions that achieve high accuracy at surprisingly low FLOPs, though with more variability in performance among candidate architectures. In all cases, improvements in accuracy are primarily driven by quantum FLOPs, since classical FLOPs remain largely fixed due to the constrained design of classical layers. Overall, these results demonstrate that FLOPs-aware HQNN search consistently identifies architectures that balance accuracy and computational efficiency. The Pareto fronts across datasets confirm that strong performance can be obtained without over-provisioning FLOPs, with dataset complexity dictating the threshold at which additional FLOPs cease to provide accuracy gains. The findings underscore the importance of incorporating FLOPs-awareness into HQNN search, as it not only enables scalable optimization for large datasets but also prevents wasted computational resources for simpler ones.